# Nes2Net: A Lightweight Nested Architecture for Foundation Model Driven Speech Anti-spoofing

### v1

@@ IV-A Datasets
We use five datasets across various scenarios, including singing voice deepfake, fully spoofed speech, adversarial attacks, and partially spoofed speech to evaluate the performance of the proposed model. Deepfake singing voice detection (SVDD) is a growing area of interest in the research community [52, 46, 53]. Accordingly, we use the Controlled Singing Voice Deepfake Detection (CtrSVDD) dataset [52, 46] as a representative example. Fully spoofed speech is the most studied category. In this work, we include two types of datasets: (1) the ASVspoof series, which comprises ASVspoof 2019 [47], ASVspoof 2021 Logical Access (LA), ASVspoof 2021 Deepfake (DF) [48], and ASVspoof 5 [49] datasets; and (2) the In-the-Wild dataset [50], which reflects real-world usage scenarios. Partially spoofed speech alters only part of an utterance to convey deceptive meaning. This emerging challenge has attracted growing attention. We use the PartialSpoof [51] dataset as a representative benchmark. Table II summarizes the datasets used in this study. Models are trained on the training set and validated on the validation set to select the best checkpoint for testing.
For CtrSVDD [46], we report results on two official test protocols, depending on whether ACESinger bona fide samples are included. A14 is excluded following the dataset’s official guidelines [46]. ASVspoof 2019 [47] is used only for training and validation, while the In-the-Wild [50], ASVSpoof 2021 LA and DF [48] datasets are used only for testing. For the recently released ASVspoof 5 dataset [49], we use its train, development, and evaluation partitions for model training, validation, and testing, respectively. For PartialSpoof [51], we follow the standard partitioning into train, development, and evaluation sets.
@@ IV-B Training Strategies
Figure 2: The cyclic learning rate schedule using cosine annealing, with a minimum value of $1\times 10^{-9}$ and a maximum value of $1\times 10^{-6}$.
TABLE III: Performance in EER (%) on the CtrSVDD evaluation set [46] with WavLM [3] front-end. Results are shown as ‘best (mean)’ over 3 runs. Params. and MMACs refer to number of parameters and million multiply-accumulate operations, respectively. w/o and w/ ACE B.F. refer to ‘without’ and ‘with’ ACESinger bona fide samples, respectively. Attack-specific EERs are computed under the ‘w/o ACE B.F’ condition. Best results are in bold; second-best are underlined.
@@
AASIST  [16] 447k 707.65 - - - - - - 2.70
AASIST Light (C=24) 159k 91.35 1.27 (1.37) 0.87 (1.00) 5.44 (5.86) 4.84 (5.65) 0.98 (1.05) 3.95 (4.35) 3.41 (3.77)
AASIST Standard(C=32) 447k 707.65 1.18 (1.28) 0.73 (0.86) 3.63 (3.86) 5.65 (5.77) 0.88 (1.00) 3.30 (3.36) 2.79 (2.89)
AASIST Large(C=40) 662k 1,091.28 1.32 (1.37) 0.87 (0.97) 3.70 (3.96) 5.04 (5.63) 0.96 (1.06) 3.19 (3.36) 2.71 (2.94)
AASIST XL(C=48) 835k 1,555.56 1.23 (1.36) 0.76 (0.92) 3.40 (4.64) 4.93 (5.55) 0.89 (1.06) 3.12 (3.62) 2.76 (3.18)
AASIST XXL(C=56) 1,087k 2,104.57 0.96 (1.20) 0.66 (0.84) 3.86 (4.15) 4.83 (5.43) 0.75 (0.95) 3.05 (3.43) 2.65 (2.95)
ResNet 611k 70.62 1.18 (1.21) 0.80 (0.93) 3.97 (5.06) 4.60 (4.86) 0.96 (1.03) 3.11 (3.61) 2.74 (3.17)
Res2Net 452k 64.93 1.26 (1.37) 0.83 (0.86) 3.59 (4.08) 4.45 (4.80) 1.08 (1.09) 3.02 (3.24) 2.61 (2.78)
ECAPA-TDNN (C=128) 497k 80.21 1.18 (1.39) 0.67 (0.85) 4.47 (5.84) 4.63 (4.96) 0.87 (1.04) 3.19 (3.74) 2.79 (3.30)
@@
As shown in Fig. 2, over 75 training epochs, we select checkpoints from the epoch with the minimum learning rate, as well as its preceding and following epochs, for validation. The best validation result is then used for testing. We use binary focal loss, a generalization of binary cross-entropy loss, with a focusing parameter ($\gamma$) of 2 and a positive class weight ($\alpha$) of 0.25. To standardize input length, each sample is randomly cropped or padded to 4 seconds during training. We adopt the Rawboost ‘parallel: (1)+(2)’ data augmentation strategy [54], as explored in [16]. WavLM is used as the front-end model for this dataset. The pre-trained and implementation of WavLM are obtained from S3PRL [^2] https://github.com/s3prl/s3prl.
ASVspoof 2019 & 2021: For the models trained on the ASVspoof 2019 [47] dataset, we follow the baseline system proposed in [15] [^3] https://github.com/TakHemlata/SSL_Anti-spoofing. Audio data are cropped or concatenated to create segments of approximately 4 seconds in duration (64,600 samples). We use the Adam optimizer [55] with a weight decay of $1\times 10^{-4}$. To reproduce the AASIST baseline [15], we reduce the original batch size from 14 to 8 due to GPU memory constraints, and halve the learning rate from $1\times 10^{-6}$ to $5\times 10^{-7}$. For Nes2Net, benefiting from its lower GPU memory consumption, we use a batch size of 12 with a learning rate of $2.5\times 10^{-7}$. The loss function used is weighted Cross Entropy. Following [15], we apply Rawboost augmentations [54], specifically ‘series: (1+2+3)’ (Algo4) and ‘series: (1+2)’ (Algo5), for AASIST baselines. For the proposed Nes2Net-X, only the former augmentation is applied. All models are trained for 100 epochs and the best checkpoint on the validation set is used for testing on the ASVspoof 2021 [48] and In-the-Wild [50] datasets.
ASVspoof 5: Both our baseline AASIST and the proposed Nes2Net-X models are trained using settings similar to those used for AASIST in the ASVspoof 2019 corpus. However, several differences apply. The final learning rate is set to $1\times 10^{-7}$, we apply data augmentation using MUSAN [56] and RIR [57], and training is stopped if there is no improvement on the development set for 5 consecutive epochs. Following the ASVspoof 5 challenge guidelines [49], we use WavLM [3] as the front-end, replacing wav2vec 2.0 [1] used in our ASVspoof 2019 experiments.
@@ V Results and Analysis
In this section, we provide an extensive validation of the proposed Nes2Net across a wide range of datasets. We begin by exploring the design motivation and development path of Nes2Net using the CtrSVDD dataset. These insights are further validated on ASVspoof 2021, ASVspoof 5, In-the-Wild, and PartialSpoof datasets. Our experimental setup covers a broad spectrum of scenarios to ensure comprehensive evaluation, including singing voice deepfakes, fully spoofed speech, adversarial attacks, partially spoofed speech, and real-world deepfake cases.
@@ V-A Studies on the CtrSVDD dataset
We conduct experiments on the CtrSVDD dataset [46], following two testing protocols: one including ACESinger bona fide samples and the other excluding them [38]. While results for both protocols are reported in Table III, our primary analysis focuses on the scenario ‘without ACESinger bona fide (w/o ACE. B.F.)’, as recommended by the dataset creators.
First, we reproduce the SOTA baseline system, AASIST Standard (C=32), achieving an EER of 2.79%, closely matching the originally reported 2.70% [16]. Under the ‘w/o ACE B.F.’ condition, the best run achieves an EER of 3.30%, with an average of 3.36% across three runs. Further experiments show that scaling up the AASIST model does not improve EER, possibly due to parameter redundancy.
@@
Back-end Dimensionality Reduction Layer Reduced Dimension $D$ Params. MMACs Pooled EER Remarks
@@
Res2Net-woDR (b=1, s=4) $\times$ - 861k 119.15 4.15 3.62 a remove dimensionality reduction layer and increase scale $s$
@@
We additionally evaluate several widely-used baseline systems, including ResNet [19], Res2Net [18], and ECAPA-TDNN [34]. ECAPA-TDNN and ResNet achieve EERs of 3.74% and 3.61%, respectively, which are slightly worse than that of AASIST. In contrast, Res2Net benefits from the advantages of multi-scale feature extraction, delivering the best average performance among the baseline systems with an EER of 3.24%. Our proposed Nes2Net outperforms all baseline systems, achieving an EER of 2.55% with the lowest computational cost. Furthermore, the enhanced version, Nes2Net-X, further improves the performance to 2.51% EER, marking the best single-model performance reported to date. Compared to Res2Net, ResNet, ECAPA-TDNN, and SOTA AASIST ($C=32$), Nes2Net-X achieves EER reductions of 23%, 30%, 33%, and 25%, respectively.
We also analyze performance across different synthetic attack types using the ‘w/o ACE B.F.’ protocol. Except for the A12 track, our model consistently achieves either the best or second-best performance, demonstrating strong generalization and robustness. Notably, the A12 track—based on Singing Voice Synthesis (SVS)—proves particularly challenging, showing higher EER across all models and highlighting a potential area for future improvement.
We observe that performance trends are consistent across both conditions—with and without ACESinger bona fide samples. Moreover, the EER is lower when ACESinger bona fide samples are included. This indicates that, even though ACESinger bona fide samples are considered out-of-domain, the trained models exhibit strong generalization capabilities and are able to classify these samples accurately.
@@ V-B The Roadmap of the Nes2Net
In this section, we introduce the roadmap from Res2Net to the proposed Nes2Net, with detailed results summarized in Table IV. To aid interpretation, we visualize the number of parameters, MACs, and EER. These are represented in Fig. 3 by circle size, the horizontal axis, and the vertical axis, respectively. In the following, we provide detailed analyses:
Investigating Res2Net: Among the baselines in Table III, the Res2Net-based back-end outperforms ResNet, AASIST, and ECAPA-TDNN on the CtrSVDD dataset. Therefore, we select it as the reference baseline for further investigation. First, we experiment with adjusting the scale $s$ of Res2Net. We observe that as $s$ increases, the number of split groups increases linearly; however, the performance shows no significant improvement (depicted as the teal blue line in Fig. 3). This may be because adding too many split groups dilutes the feature representation, leading to redundancy and limited benefit.
Next, we explore varying the dimensionality of the output features from the DR layer (referred to as Reduced Dimension $D$, depicted as the steel gray line in Fig. 3). Reducing $D$ to 64 significantly lowers model size and MACs compared to the default $D=128$, but leads to substantial performance degradation, increasing EER from 3.02% to 4.32%. Conversely, increasing $D$ to 256 results in a much larger model size and MACs but still leads to worse performance than $D=128$. This may be because a larger $D$ introduces over-parameterization and noise. This likely explains why $D=128$ is commonly adopted in SOTA models [15, 16].
@@
We further evaluate the performance of ResNet-woDR with different scales $s$ (depicted as the green line in Fig. 3). The best performance is observed with $s=64$, achieving an EER of 2.73%, which surpasses the best Res2Net baseline. Increasing $s$ beyond this point leads to a decline in performance, likely due to the following factors:
@@
Based on the optimal $s$, we increase the number of blocks $b$ to deepen the model (depicted as the light pink line in Fig. 3). However, no further performance improvement is observed. This may be because the deeper architecture does not effectively leverage the additional parameters, resulting in diminishing performance gains. It may also increase the risk of overfitting.
@@
The Novel Nested Design: Prior experiments demonstrate that removing the DR layer enhances the performance of Res2Net. We believe that directly extracting information from high-dimensional speech foundation model features avoids the information loss introduced by DR. Our experiments with variations in scale, depth, and dimensionality show that an EER of 2.73% represents the performance bottleneck for this design.
@@
Compared to ResNet-woDR, the proposed Nes2Net adopts a novel nested design that enhances flexibility and significantly boosts the model’s representational capacity. Processing larger feature subsets in the outer layer facilitates better interactions across channels within each nested layer. Furthermore, the integrated local cross-channel attention mechanism enhances feature selection while mitigating redundancy, addressing limitations in prior designs. This architectural refinement overcomes the performance limitations observed in the original Res2Net design. As a result, Nes2Net and its enhanced variant Nes2Net-X surpass the earlier performance bottleneck, achieving average EERs of 2.55% and 2.51%, respectively.
V-C Activation Function
TABLE V: Performance in EER (%) on the CtrSVDD Eval set [46] for comparing with proposed Nes2Net with different activation functions. The results are reported as the format of ‘best (mean)’ across 3 runs.
Front-end Back-end Activation Function Pooled EER
w/o ACE. B.F. w/ ACE. B.F.
WavLM Nes2Net-X ReLU [60] 2.48 (2.51) 2.20 (2.24)
GELU [61] 2.52 (2.56) 2.22 (2.26)
SiLU [62] 2.52 (2.55) 2.22 (2.25)
SELU [63] 2.28 (2.69) 2.02 (2.37)
Activation functions like Rectified Linear Unit (ReLU) [60], Scaled Exponential Linear Units (SELU) [63], Gaussian Error Linear Units (GELU) [61], and Sigmoid Linear Unit (SiLU) [62] address common deep learning challenges, such as vanishing gradients. By introducing diverse characteristics like improved gradient flow, self-normalization, or smooth transitions, these functions aim to enhance model convergence, stability, and overall performance in complex scenarios.
We experiment with replacing the default ReLU activation function with several alternatives and present the results in Table V. GELU and SiLU do not yield any performance improvements over ReLU. Although SELU achieves better performance in one of the three runs, its average performance does not improve. This suggests higher variance across runs. This suggests that SELU does not offer a consistent advantage over ReLU in our experimental setup. Therefore, we retain ReLU as the default activation function in our model.
V-D Studies on the ASVspoof 2021 dataset
The ASVspoof series datasets are widely used as benchmarks for advancing research in detecting spoofed speech [47, 48]. Following standard protocol, we train our models on ASVspoof 2019 [47] and evaluate them on ASVspoof 2021 Logical Access (LA) and Deepfake (DF) tasks [48]. The LA task focuses on detecting synthetic and voice-converted speech transmitted over telephony systems, introducing challenges related to channel effects and transmission variability. In contrast, the DF task targets detecting manipulated, compressed speech data commonly found on online platforms. This reflects real-world scenarios where deepfake audio circulates, making the DF task a valuable benchmark for evaluating deepfake detection systems.

### v2

@@ IV-A Datasets
We use five datasets across various scenarios, including singing voice deepfake, fully spoofed speech, adversarial attacks, and partially spoofed speech, to evaluate the performance of the proposed model. Singing voice deepfake detection (SVDD) is a growing area of interest in the research community [52, 46, 53]. The CtrSVDD dataset [52, 46] offers structured attack types and official evaluation protocols, making it suitable for systematic architecture exploration. As a newly collected resource, it captures recent spoofing techniques, providing a more challenging and relevant benchmark for modern anti-spoofing systems. We therefore adopt it as a representative example. Moreover, fully spoofed speech is the most studied category. In this work, we include two categories of datasets: (1) the ASVspoof series, which comprises ASVspoof 2019 [47], ASVspoof 2021 Logical Access (LA), ASVspoof 2021 Deepfake (DF) [48], and ASVspoof 5 [49]; and (2) the In-the-Wild dataset [50], which reflects real-world usage scenarios. Partially spoofed speech alters only part of an utterance to convey deceptive meaning. This emerging challenge has attracted growing attention. We use the PartialSpoof [51] dataset as a representative benchmark. Table II summarizes the datasets used in this study. Models are trained on the training set and validated on the validation set to select the best checkpoint for testing.
For CtrSVDD [46], we report results on two official test protocols, according to whether ACESinger bona fide samples are included. The ‘A14’ attack type of the CtrSVDD dataset is excluded following the official guidelines [46]. ASVspoof 2019 [47] is used only for training and validation, while the In-the-Wild [50], ASVspoof 2021 LA and DF [48] datasets are used only for testing. For the recently released ASVspoof 5 dataset [49], we use its train, development, and evaluation partitions for model training, validation, and testing, respectively. For PartialSpoof [51], we follow the standard partitioning into train, development, and evaluation sets.
@@ IV-B Training Strategies
Figure 2: The cyclic learning rate schedule using cosine annealing.
@@
[…]
@@
AASIST (C=32) [16] 447k 707.65 - - - - - - 2.70
AASIST Light (C=24) $\dagger$} 159k 91.35 1.27 (1.37) 0.87 (1.00) 5.44 (5.86) 4.84 (5.65) 0.98 (1.05) 3.95 (4.35) 3.41 (3.77)
AASIST Standard(C=32) $\dagger$} 447k 707.65 1.18 (1.28) 0.73 (0.86) 3.63 (3.86) 5.65 (5.77) 0.88 (1.00) 3.30 (3.36) 2.79 (2.89)
AASIST Large(C=40) $\dagger$} 662k 1,091.28 1.32 (1.37) 0.87 (0.97) 3.70 (3.96) 5.04 (5.63) 0.96 (1.06) 3.19 (3.36) 2.71 (2.94)
AASIST XL(C=48) $\dagger$} 835k 1,555.56 1.23 (1.36) 0.76 (0.92) 3.40 (4.64) 4.93 (5.55) 0.89 (1.06) 3.12 (3.62) 2.76 (3.18)
AASIST XXL(C=56) $\dagger$} 1,087k 2,104.57 0.96 (1.20) 0.66 (0.84) 3.86 (4.15) 4.83 (5.43) 0.75 (0.95) 3.05 (3.43) 2.65 (2.95)
ResNet $\dagger$} 611k 70.62 1.18 (1.21) 0.80 (0.93) 3.97 (5.06) 4.60 (4.86) 0.96 (1.03) 3.11 (3.61) 2.74 (3.17)
Res2Net $\dagger$} 452k 64.93 1.26 (1.37) 0.83 (0.86) 3.59 (4.08) 4.45 (4.80) 1.08 (1.09) 3.02 (3.24) 2.61 (2.78)
ECAPA-TDNN (C=128) $\dagger$} 497k 80.21 1.18 (1.39) 0.67 (0.85) 4.47 (5.84) 4.63 (4.96) 0.87 (1.04) 3.19 (3.74) 2.79 (3.30)
@@
※: XWSB is an ensemble-like model that combine two SSL front-ends [39], while all other models in Table III are based on single SSL front-end.
As shown in Fig. 2, over 75 training epochs, we select checkpoints from the epoch with the minimum learning rate, as well as its preceding and following epochs, for validation. The best validation result is then used for testing. We use binary focal loss [54], a generalization of binary cross-entropy loss, with a focusing parameter ($\gamma$) of 2 and a positive class weight ($\alpha$) of 0.25. To standardize input length, each sample is randomly cropped or padded to 4 seconds during training. We adopt the Rawboost ‘parallel: (1)+(2)’ data augmentation strategy [55], as explored in [16]. WavLM is used as the front-end model for this dataset. The pre-trained and implementation of WavLM are obtained from S3PRL [^2] https://github.com/s3prl/s3prl.
ASVspoof 2019 & 2021: For the models trained on the ASVspoof 2019 [47] dataset, we follow the baseline system proposed in [15] [^3] https://github.com/TakHemlata/SSL_Anti-spoofing. Audio data are cropped or concatenated to create segments of approximately 4 seconds in duration (64,600 samples) for both training and testing. We use the Adam optimizer [56] with a weight decay of $1\times 10^{-4}$. To reproduce the AASIST baseline [15], we reduce the original batch size from 14 to 8 due to GPU memory constraints, and halve the learning rate from $1\times 10^{-6}$ to $5\times 10^{-7}$. For Nes2Net, benefiting from its lower GPU memory consumption, we use a batch size of 12 with a learning rate of $2.5\times 10^{-7}$. The loss function used is weighted Cross Entropy. Following [15], we apply Rawboost augmentations [55], specifically ‘series: (1+2+3)’ (Algo4) and ‘series: (1+2)’ (Algo5), for AASIST baselines. For the proposed Nes2Net-X, only the former augmentation is applied. All models are trained for 100 epochs and the best checkpoint on the validation set is used for testing on the ASVspoof 2021 [48] and In-the-Wild [50] datasets.
ASVspoof 5: Both our AASIST baseline and the proposed Nes2Net-X models are trained using settings similar to those used for AASIST in the ASVspoof 2019 corpus. However, several differences apply. The final learning rate is set to $1\times 10^{-7}$, we apply data augmentation using MUSAN [57] and RIR [58], and training is stopped if there is no improvement on the development set for 5 consecutive epochs.
@@ V Results and Analysis
All Equal Error Rate (EER) results in this work are reported as ‘best (mean)’ over multiple runs. For cited results that (1) are based on a single run, (2) report only the best result, or (3) lack sufficient details, only a single value is presented.
@@ V-A Studies on the CtrSVDD dataset
We conduct experiments on the CtrSVDD dataset [46], following two testing protocols: one including ACESinger bona fide samples and the other excluding them [38]. While results for both protocols are reported in Table III, our primary analysis focuses on the scenario ‘without ACESinger bona fide (w/o ACE. B.F.)’, as recommended by the dataset creators. Since AASIST (C=32) in our prior work [16], as well as SLS and XWSB [39], were evaluated during the CtrSVDD Challenge 2024, portions of their test sets differ from the current official protocol. As a result, the EER by attack type is not directly comparable. To ensure a fair comparison, we re-implemented the AASIST (C=32) system under the official protocol and used it as our baseline, referred to as AASIST Standard (C=32) in Table III, achieving an EER of 2.79%, which is close to the originally reported 2.70% [16]. Under the ‘w/o ACE B.F.’ condition, the best run achieves an EER of 3.30%, with an average of 3.36% across three runs. Further experiments show that scaling up the AASIST model does not improve mean EER, possibly due to parameter redundancy.
@@
Back-end
Dimensionality
Reduction Layer
Reduced
Dimension $D$
Params. MMACs Pooled EER Remarks
@@
Res2Net-woDR (b=1, s=4) $\times$ - 861k 119.15 4.15 3.62
a remove dimensionality
reduction layer
and increase scale $s$
@@
We additionally evaluate several widely-used baseline systems, including ResNet [19], Res2Net [18], and ECAPA-TDNN [34]. ECAPA-TDNN and ResNet achieve EERs of 3.74% and 3.61%, respectively, which are slightly worse than that of AASIST. In contrast, Res2Net benefits from the advantages of multi-scale feature extraction, delivering the best average performance among the baseline systems with an EER of 3.24%. Our proposed Nes2Net outperforms all baseline systems, achieving a mean EER of 2.55% with the lowest computational cost. Furthermore, the enhanced version, Nes2Net-X, further improves the performance to 2.51% EER, marking the best single-model performance reported to date. Compared to Res2Net, ResNet, ECAPA-TDNN, and SOTA AASIST ($C=32$), Nes2Net-X achieves EER reductions of 23%, 30%, 33%, and 25%, respectively.
We also analyze performance across different synthetic attack types using the ‘w/o ACE B.F.’ protocol. Except for the ‘A12’  attack type [46], our model consistently achieves either the best or second-best performance, demonstrating strong generalization and robustness. Notably, the ‘A12’  attack type, based on Singing Voice Synthesis (SVS), proves particularly challenging, showing higher EER across all models and highlighting a potential area for future improvement.
We observe that performance trends are consistent across both conditions, with and without ACESinger bona fide samples. Moreover, the EER is lower when ACESinger bona fide samples are included. This indicates that, even though ACESinger bona fide samples are considered out-of-domain, the trained models exhibit strong generalization capabilities and are able to classify these samples accurately.
@@ V-B The Roadmap of the Nes2Net
In this section, we introduce the roadmap from Res2Net to the proposed Nes2Net, with detailed results summarized in Table IV. All systems are implemented and evaluated under a unified framework for fair comparison. To aid interpretation, we visualize the number of parameters, MACs, and EER. These are represented in Fig. 3 by circle size, the horizontal axis, and the vertical axis, respectively. In the following, we provide detailed analyses:
Investigating Res2Net: Among the baselines in Table III, the Res2Net-based back-end outperforms ResNet, AASIST, and ECAPA-TDNN on the CtrSVDD dataset. Therefore, we select it as the reference baseline for further investigation. First, we experiment with adjusting the scale $s$ of Res2Net. We observe that as $s$ increases, the number of split groups increases linearly; however, the performance shows no significant improvement (depicted as the teal blue line in Fig. 3). This may be because adding too many split groups dilutes the feature representation, leading to redundancy.
Next, we explore varying the dimensionality of the output features from the DR layer (referred to as Reduced Dimension $D$, depicted as the steel gray line in Fig. 3). Reducing $D$ to 64 significantly lowers model size and MACs, compared to the default $D=128$, but leads to substantial performance degradation, increasing EER from 3.02% to 4.32%. Conversely, increasing $D$ to 256 results in a much larger model size and MACs but still leads to worse performance than $D=128$. This may be because a larger $D$ introduces over-parameterization and noise. This may explain why $D=128$ is commonly adopted in SOTA models [15, 16].
@@
We further evaluate the performance of ResNet-woDR with different scales $s$ (depicted as the green line in Fig. 3). The best performance is observed with $s=64$, achieving a mean EER of 2.97%, which surpasses the best Res2Net baseline. Increasing $s$ beyond this point leads to a decline in performance, likely due to the following factors:
@@
Based on the optimal $s$, we increase the number of blocks $b$ to deepen the model (depicted as the light pink line in Fig. 3). However, no further performance improvement is observed. This could be attributed to the deeper architecture’s limited ability to effectively utilize the additional parameters, resulting in diminishing performance gains. It may also increase the risk of overfitting.
@@
The Novel Nested Design: Prior experiments demonstrate that removing the DR layer enhances the performance of Res2Net. We believe that directly extracting information from high-dimensional speech foundation model features avoids the information loss introduced by DR. Our experiments with variations in scale, depth, and dimensionality show that a mean EER of 2.97% marks a performance bottleneck for this design.
@@
Compared to ResNet-woDR, the proposed Nes2Net adopts a novel nested design that enhances flexibility and significantly boosts the model’s representational capacity. Processing larger feature subsets in the outer layer facilitates better interactions across channels within each nested layer. Furthermore, the integrated local cross-channel attention mechanism enhances feature selection while mitigating redundancy, addressing limitations in prior designs. This architectural refinement overcomes the performance limitations observed in the original Res2Net design. As a result, Nes2Net and its enhanced variant Nes2Net-X surpass the earlier performance bottleneck, achieving mean EERs of 2.55% and 2.51%, respectively.
V-C Studies on the ASVspoof 2021 dataset